OpenStudy (anonymous):

find all values of x to the nearest tenth 2secx+10=0 Help please

OpenStudy (anonymous):

all values?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[2\sec(x)+10=0\\ 2\sec(x)=-10\\ \sec(x)=-5\\ \cos(x)=-\frac{1}{5}\\ \cos(x)=-.2\] after that you are on your own

OpenStudy (anonymous):

what do you do afterwards?

OpenStudy (anonymous):

really need help please

OpenStudy (anonymous):

take the arccosine with a calculator that is all i can do

OpenStudy (anonymous):

but i think there is more than one answer right?

OpenStudy (anonymous):

since it is negative

OpenStudy (anonymous):

hey satellite can u fan me so that i can send you some messages where you can possible help me ?

OpenStudy (anonymous):

there are an infinite number of answers are you working in degrees or radians?

OpenStudy (anonymous):

between 0 degrees and 360 degrees sorry

OpenStudy (anonymous):

ok one answer is this http://www.wolframalpha.com/input/?i=arccos%28-.2%29

OpenStudy (anonymous):

looks like \(101.5^\circ\)

OpenStudy (anonymous):



OpenStudy (anonymous):

other answer is \[360^\circ-101.5^\circ\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw